OpenStudy (anonymous):

Given the points (-2, 1) and (5, -5) b) Write an equation in point-slope form of the line through the points. c) Convert the equation to slope-intercept form d) Convert the equation to standard form. e) Graph the equation. You may use the axes provided, or create your own graph.

OpenStudy (anonymous):

find slope using m= (y2-y1) / (x2-x1) where y2 and x2 are (5,-5) and y1 and x1 are (-2,1) then use y-y1= m * ( x-x1) or y-y2=m*(x-x2) to find your equation

OpenStudy (anonymous):

doing that you will get the answer to question 2 automatically

OpenStudy (anonymous):

im still a little bit lost